The management of an automobile company tries to find out the effectiveness of staff salary increment on performance of staff pls step by step.. and the word limt for interperation (100-125 words)
Answer #1

Q1) a)

Find X · Y and X2 as it was done in the table below. X Y XY X X 1 10 10 13 100 169 1.5 19.5 1.9 15 17 28.5 40.8 2.4 2.7 225 2

here your alpha is 'a' which is 1.315 and beta is 'b' which is 0.057.

А B с D E 1 2 3 41.67% 4 5 6 7 oc salary increment performance 10 1 performance below average 13 1.5 15 1.9 17 2.4 19 2.7 23

from above 41.67% of staff perfomed below average. You can see red highlighted cells those are the numbers below average=2.992. so percentage = (5/12)*100 = 41.67%

average increment is the mean of increment column which is 29.417.

Interpretation-

as you can see regression equation is

y=1.315+0.057*x so a postive slope coefficient b shows that if we increase the salary of employee their perfromance increases that's pne thing. And quantitatively if we increase the salry by 1% the performace increases by 0.057 % which means let's say someone gets an increment of 30% so his performance is expected to increase by 1.71%.

but don't think that if we increase salary by say 1000% then perfromance will aslo increase by 0.057*1000= 57%

because human capability is limited this model will give results for a short period of time which means for some values of increments around the data.

Q2)

a) correlation is required when we want to have a quick galnce at if a linear relationship exists between two variables and if it exists what is the strength and direction of the relationship.

for eg in the question above the owner of the company is interested in knowing if expenditure on advertisement increases sales or not.

b)

Find X Y , X2 and Y2 as it was done in the table below. Υ.Υ X 7450 Y 89 7921 5230 39 1521 77 5929 6250 5975 52 X Y 663050 203

c)

correlation coefficient always lies between +1 and -1. r = 0.9427 is the correlation coefficient for our question which shows a very strong positive correlation as it is very close to 1 . It means yes there exists a strong linear and direct relationship between amount of expenditure on advertisemnet and number of new customers.
So if we increase advertisement we can expect to see a increase in number of new customers.

I hope this helps you

please upvote

Thanks